To make a solution, 40 g of a nonelectrolyte is dissolved into 300 g of water. The freezing point depression of the solution is measured to be -7°C. Calculate the molar mass of the solute.


1
Expert's answer
2022-03-20T06:38:09-0400

The lowering (depression) of the freezing point of the solvent can be represented using the following equation: 

Δt = i × Kf × m

where:

Δt = the change in freezing point

i = van't Hoff factor

Kf = the freezing point depression constant

m = the molality of the solute


The normal freezing point of pure water is 0°C

Therefore, Δt = 0°C − (−7°C) = 7°C

i = 1 (for nonelectrolyte)

Kf = 1.86 °C kg mol¯1 (for water)


Thus,

m = Δt / (i × Kf)

Molality of solute = (7°C) / (1 × 1.86 °C kg mol¯1) = 3.7634 mol kg¯1


Molality of solute = Moles of solute / Kilograms of solvent

Kilograms of solvent = Kilograms of water = (300 g) × (1 kg / 1000 g) = 0.3 kg

Therefore,

Moles of solute = Molality of solute × Kilograms of water

Moles of solute = (3.7634 mol kg¯1) × (0.3 kg) = 1.129 mol


Moles = Mass / Molar mass

Therefore,

Molar mass of solute = Mass of solute / Moles of solute

Molar mass of solute = (40 g) / (1.129 mol) = 35.43 g/mol

Molar mass of solute = 35.43 g/mol


Answer: The molar mass of the solute is 35.43 g/mol